let p=0.09q + 50 be the supply equation for a manufacturer. the demand equation for his product is p = 0.07q + 65. if a tax of n1.50 per unit is to be imposed on the manufacturer, how will the original equilibyium price be effected if the demand remains the same?
Answer by JoelSchwartz(130)

The equilibrium price is when the supply equation is equal to the demand equation.
.09q+50=.07q+65
.02q=15
100*.02q=100*15
2q=1,500
q=750
the equilibrium price is .07*750+65=117.5
when i add the tax of 1.5 per unit to the equilibrium price i get 1.57*750+65=1,242.5